# Scheduling Lattice Surgery with Magic State Cultivation

### v1

@@ I Introduction
The immense potential of quantum computing is tempered by the high error rates observed in existing hardware. As a result, scalable architectures must rely on Quantum Error Correction Codes (QECCs), which encode fewer logical qubits into many physical qubits to detect and correct errors. The surface code [kitaev2003fault, fowler2012surface] is a widely studied QECC because it requires only nearest-neighbor interactions on a two-dimensional grid, making it especially well-suited to hardware that provides planar connectivity between physical qubits (e.g. superconducting qubits). Long-range interactions between logical qubits are provided via lattice surgery, the dynamic merging and splitting of code patches using targeted measurements and hardware manipulations [horsman2012surface].
@@
Surface codes face a fundamental challenge: they cannot directly implement arbitrary quantum gates. Instead, they rely on a restricted set of operations supplemented by magic states, which are specially prepared quantum states that enable universal computation [bravyi2005universal]. Magic state preparation has historically required distillation protocols that consume thousands of noisy auxiliary states to produce a single high-fidelity magic state, creating enormous resource overhead [fowler2018low].
@@
The execution of a high-level quantum circuit on a target QECC such as a surface code employs a compiler to transform a sequence of single- and two-qubit interactions into multi-qubit Pauli product measurements for Pauli Basis Computation [bravyi2016pbc, Litinski_2019]. The goal of the compiler is to minimize the space (qubits) and time (logical cycles) of the compiled program. Doing so efficiently requires managing the additional logical qubits needed for routing long-distance operations and preparing magic state. To address this challenge, existing surface-code architectures assign fixed roles to different regions of the grid: data qubits store program state, bus qubits enable routing via lattice surgery, and dedicated magic patches produce magic states. These designs typically assume large, space-intensive magic-state distillation factories [bravyi2005universal] placed statically along the grid boundary.
@@
As an alternative to distillation, recent work has introduced magic-state cultivation, a more resource-efficient method that prepares a magic state using the footprint of only a single logical qubit [gidney2024cultivation]. Using cultivation, magic state production can be assigned to any qubit, not just those on the grid boundary. This enables a new approach, which we call Pure Magic: a dynamic lattice surgery scheduling approach that repurposes qubits on demand. In this approach, there are only only data and ancilla qubits. Ancillae continuously cultivate magic states, and when an ancilla qubit is needed for routing, cultivation is canceled. After routing, the cultivation process restarts immediately.
@@
We implement the Pure Magic approach through a new scheduler, which determines the order in which to execute a sequence of Pauli products comprising a circuit. Many of these Pauli products can be simultaneously executed on the surface code, given adequate routing and magic state resources. Determining how to order the execution to maximize parallelism is the Lattice Surgery Scheduling Problem (LSSP), which can be solved by a greedy Steiner tree packing algorithm [silva2024lssp].
@@
We evaluate the Pure Magic scheduler on benchmarks from Benchpress [nation2025benchmarking], augmented with randomly generated circuits. Across these tests, Pure Magic reduces scheduling volume—a proxy for execution time and logical error—by 19% to 220% relative to state-of-the-art techniques. By re-purposing ancilla patches for both routing and cultivation, Pure Magic eliminates dedicated bus qubits and significantly increases the number of active cultivators. Moreover, the dynamic reuse of ancilla patches mitigates the long-tail behavior of probabilistic cultivation: routing frequently interrupts the slowest cultivation attempts. This leads to an effective $2.4\times$-$6.7\times$ reduction in average cultivation time. These advantages are most pronounced in highly parallel circuits with substantial demand for magic states. Reducing average magic-state generation time is particularly valuable because the overall speed of a fault-tolerant quantum computer is governed by how quickly magic states can be produced and consumed.
@@ II-A Surface Codes
We model surface-code logical qubits as patches arranged on a two-dimensional grid [horsman2012surface]. This abstraction ensures that all required physical operations map cleanly onto a planar architecture. The edges of each patch play a central role: merging patches along specific edges performs joint parity measurements of the corresponding logical operators [horsman2012surface, fowler2018low]. Figure 2(b) shows a single surface-code patch with its $X$- and $Z$-type edges labeled. In our architecture, we employ double-qubit patches, represented as $2\times 1$ rectangles. Each such patch encodes two logical qubits within a single surface-code region. This layout places the $X$ and $Z$ operators for each qubit along the same boundary of the rectangle, enabling more efficient multi-qubit interactions and simplifying lattice-surgery operations (Figure 2(c)).
@@ II-B Pauli Basis Computation via Lattice Surgery
Input circuits are assumed to use the standard Clifford+T gate set [fowler2012surface, Litinski_2019, O_Gorman_2017]. Our compiler transforms Clifford+T circuits into sequences of Pauli product rotations by rewriting gates into their Pauli rotation equivalents (e.g., $S=Z_{\pi/4}$, $T=Z_{\pi/8}$). All Clifford operations are commuted to the end of the circuit and absorbed into measurements using the Clifford tableau method [aaronson_2004, silva2024lssp].
@@
Lattice surgery [horsman2012surface] provides an efficient method to execute Pauli products through merge and split operations, as shown in Figure 3. These form the native instruction set of lattice surgery architectures.
@@ II-C Magic State Cultivation
Universal quantum computation requires magic states, which cannot be initialized fault-tolerantly. Traditional approaches rely on distillation, a process that consumes thousands of noisy resource states to produce a single high-fidelity one, leading to extremely large overheads [bravyi2005universal, fowler2012surface].
Recent work by Gidney et al. [gidney2024cultivation] introduced cultivation, a far more efficient approach using the same physical resources as a single logical qubit. Cultivation proceeds in three stages. First, injection creates the initial encoded T state in a low-distance code. Second, cultivation gradually increases reliability using “check-grow-stabilize” cycles with post-selection (shots are discarded when any error is detected). Finally, the escape stage quickly boosts the code distance to prevent the highly reliable state from degrading. At this point the code is “grafted” into a matchable surface-code-like patch. Cultivation’s efficiency enables more practical algorithm implementations and compact architectures [gidney2025factor2048bitrsa], but existing scheduling algorithms have not thoroughly studied its implications for resource management.
@@ II-D Other approaches to FTQC with Surface Codes
Prior work on lattice-surgery compilation spans several distinct approaches. LaSsynth [Tan_2024] introduces a SAT encoding of lattice surgery that can produce depth-optimal implementations of small-scale primitives. Although optimal, the method does not scale to the compilation of full application-level circuits. Building on SAT-based formulations, DASCOT [molavi_2025] presents an encoding of the Lattice Surgery Scheduling Problem together with heuristic approximations to optimal solutions; however, it assumes that magic states are always available when needed and operates directly on Clifford+T circuits rather than translating them into Pauli products. Similarly, liblsqecc [leblond_2024] proposes a framework for directly executing Clifford+T gates but considers only earlier generations of magic-state distillation and does not incorporate recent cultivation techniques.
Architectural work has also explored ways of mitigating resource bottlenecks. LSQCA [Kobori_2025] introduces a design in which surface-code patches expected to be used soon are kept in sparse compute regions, while others are stored in denser regions; the movement between these regions helps conceal the latency of magic-state generation.
@@ III Scheduling
The goal of the scheduler is to solve the Lattice Surgery Scheduling Problem (LSSP). Given a circuit comprised of a sequence of Pauli products with $L$ qubits, the scheduling algorithm must compute an ordering of the Pauli products that minimizes volume, thereby minimizing the probability of error. The scheduler must also determine the grid layout of the logical qubits, with $L$ data qubits and additional ancilla qubits for routing and magic injection, for a total of $N$ qubits with $N>L$.
@@
Here $\epsilon$ is the error rate per physical qubit per cycle, $\epsilon_{th}$ is the error threshold, and $d$ is the code distance. Equation 1 assumes that physical errors (e.g. bit-flips, phase flips) are independent and identically distributed, both in time and space. In practice the error rate can be much more complex with dependent effects, governed by the underlying hardware. We follow other analyses in this area [fowler2012surface, raussendorf2007fault] and use the simple error model. Minimizing volume minimizes errors under this error model.
@@ III-B Magic Injection Using Cultivation
We conducted Markov Chain Monte Carlo (MCMC) simulations using the survivorship probabilities of each stage of magic state cultivation outlined in Figure 15 of Gidney et. al. [gidney2024cultivation]. We simulated the cultivation of fault-distance-3 magic states so as to attain an error probability of $p_{L}\approx 4\times 10^{-6}$. Fitting the simulation data to Equation 3 yields a value of $\lambda=0.00227$. The duration of cultivation in scheduling cycles can be computed by taking samples from this exponential distribution then dividing by the code distance. Figure 7 shows the cumulative distribution function of $10^{6}$ simulated runs of cultivation.
The relationship between noise model, intermediate fault-distance, and final code distance is highly complex. Furthermore, recent work has demonstrated how cultivation can be made more time efficient  [vaknin2025efficientmagicstatecultivation, sahay2025foldtransversalsurfacecodecultivation]. For this reason, we match the parameters presented in Figure 15a of Gidney at. al. [gidney2024cultivation], but leave $\lambda$ as a controllable hyperparameter of our simulation framework.
@@ III-C Packing Steiner Forests
The scheduling algorithm attempts to maximize the number of non-overlapping, independent Pauli products scheduled in each cycle. As described previously, each Pauli product is scheduled as a tree on the grid. Ideally this is a Steiner tree, which connects all the terminals (the set of leaves representing the Pauli product operators) via the shortest paths. However, finding a Steiner tree is in general an NP-hard problem [lin2002terminal] so we find approximate Steiner trees using a variant of the shortest path heuristic [kou1981fast]. This yields a tree with weight (path length) within 22/S2-2/S of the Steiner tree, where $S$ is the number of terminals. Although there are other algorithms with improved lower bounds on the approximation, in practice the shortest path heuristic provides sufficient accuracy for the sizes of trees for typical circuits.
Maximizing the number of simultaneously scheduled Pauli products is akin to solving the Steiner forest packing problem [lau2005packing], which is also NP-hard. We find an approximate solution with a greedy MINFIT algorithm, as shown in Algorithm 1. The algorithm begins with the set of available Pauli products (those that have no parents in the task graph) and all squares in the grid. In each iteration of the while loop (line 4), for every Pauli product  an approximate Steiner tree is calculated for the available patches (line 6), and the Pauli product with the shorter path (the minimum fit) is chosen as the next to schedule (lines 7 to 10). That Pauli product is added to the Pauli products to schedule next (line 12), and removed from the list of available Pauli products (line 13), while the patches used by the tree for the Pauli product are removed from the available patches (line 14). Once all available Pauli products have been attempted, the cycle is complete. The MINFIT algorithm is repeated every cycle until no more unscheduled Pauli products remain, resulting in a schedule that indicates which Pauli products are to be executed at each cycle.
@@ IV Pure Magic Scheduling
One of the most significant performance constraints in surface-code quantum computation is the production of magic states. Even with cultivation, T-state preparation remains slow. Using the parameters reported in Gidney et. al. [gidney2024cultivation], our simulations indicate an average cultivation time of roughly 26 logical cycles. The latency of cultivation remains a large source of run time overhead, which is detrimental to performance for near-term architectures.
@@ V-A Test Circuits
We used a set of 17 benchmark circuits drawn from the Benchpress  [^2] The ”Heisenberg” benchmarks are the ”Square Heisenberg” in the Benchpress repository. quantum circuit suite [nation2025benchmarking], as shown in Table I. These circuits span many domains of interest (Quantum Machine Learning, Ising Models, Heisenberg Models, Convex Optimization). The circuits were lowered from their initial gate set to the standard Clifford+T gateset using BQSKit [bqskit]. Clifford+T circuits were then converted to a sequence of Pauli products using the algorithm described in Silva et. al. [silva2024lssp]. The layers are derived from the task graph of the resulting Pauli products.
@@ V-D Comparison with Prior Work
This work is closest in nature to that of Silva et. al. [silva2024lssp]. In their paper, they present results with a set of specially generated benchmark circuits. For our work, we chose instead to use well-known benchmarks from the Benchpress suite. However, we have run our schedulers on the Silva et. al. benchmarks for comparison, using the versions they label “transpiled”. We present our scheduler performance compared to the results given in $E(N)$ column in Table 3 from Silva et. al. Specifically, we use our bus routing scheduler only, to get as close a comparison as possible to their results.
@@ VI-B Compatibility with other codes and execution models
Although we present the Pure Magic architecture in the context of the surface code, the conceptual approach is not code specific. Any scheme in which magic states may be cultivated locally (e.g. with color codes and other codes [yoder2025tourgrossmodularquantum, steane1996qec]) could adopt a similar strategy. The key requirement is that ancillary space dedicated to routing be reallocated towards cultivation.
@@ VI-C Comparing and benchmarking QECCs
In general, QECCs are compared based on their qubit overhead, error suppression capabilities, and suitability for different hardware architectures. For example, while surface codes are robust, they are often criticized for their high physical to logical qubit encoding overhead. Whenever a code is compared against surface codes, the comparison [yoder2025tourgrossmodularquantum, viszlai2024matchinggeneralizedbicyclecodesneutral] uses static logical qubit layouts with dedicated bus and magic generation regions. This architecture has a fixed data to ancilla qubit ratio, and codes that use fewer physical qubits to encode a logical qubit tend to win against surface codes.

### v2

@@ I Introduction
The immense potential of quantum computing is tempered by the high error rates observed in existing hardware. As a result, scalable architectures must rely on Quantum Error Correction Codes (QECCs), which encode fewer logical qubits into many physical qubits to detect and correct errors. The surface code [13, 8] is a widely studied QECC because it requires only nearest-neighbor interactions on a two-dimensional grid, making it especially well-suited to hardware that provides planar connectivity between physical qubits (e.g. superconducting qubits). Long-range interactions between logical qubits are provided via lattice surgery, the dynamic merging and splitting of code patches using targeted measurements and hardware manipulations [12].
@@
Surface codes face a fundamental challenge: they cannot directly implement arbitrary quantum gates. Instead, they rely on a restricted set of operations supplemented by magic states, which are specially prepared quantum states that enable universal computation [3]. Magic state preparation has historically required distillation protocols that consume thousands of noisy auxiliary states to produce a single high-fidelity magic state, creating enormous resource overhead [7].
@@
The execution of a high-level quantum circuit on a target QECC such as a surface code employs a compiler to transform a sequence of single- and two-qubit interactions into multi-qubit Pauli product measurements for Pauli Basis Computation [4, 19]. The goal of the compiler is to minimize the space (qubits) and time (logical cycles) of the compiled program. Doing so efficiently requires managing the additional logical qubits needed for routing long-distance operations and preparing magic state. To address this challenge, existing surface-code architectures assign fixed roles to different regions of the grid: data qubits store program state, bus qubits enable routing via lattice surgery, and dedicated magic patches produce magic states. These designs typically assume large, space-intensive magic-state distillation factories [3] placed statically along the grid boundary.
@@
As an alternative to distillation, recent work has introduced magic-state cultivation, a more resource-efficient method that prepares a magic state using the footprint of only a single logical qubit [9]. Using cultivation, magic state production can be assigned to any qubit, not just those on the grid boundary. This enables a new approach, which we call Pure Magic: a dynamic lattice surgery scheduling approach that repurposes qubits on demand. In this approach, there are only only data and ancilla qubits. Ancillae continuously cultivate magic states, and when an ancilla qubit is needed for routing, cultivation is canceled. After routing, the cultivation process restarts immediately.
@@
We implement the Pure Magic approach through a new scheduler, which determines the order in which to execute a sequence of Pauli products comprising a circuit. Many of these Pauli products can be simultaneously executed on the surface code, given adequate routing and magic state resources. Determining how to order the execution to maximize parallelism is the Lattice Surgery Scheduling Problem (LSSP), which can be solved by a greedy Steiner tree packing algorithm [25].
@@
We evaluate the Pure Magic scheduler on benchmarks from Benchpress [21], augmented with randomly generated circuits. Across these tests, Pure Magic reduces scheduling volume—a proxy for execution time and logical error—by 19% to 220% relative to state-of-the-art techniques. By re-purposing ancilla patches for both routing and cultivation, Pure Magic eliminates dedicated bus qubits and significantly increases the number of active cultivators. Moreover, the dynamic reuse of ancilla patches mitigates the long-tail behavior of probabilistic cultivation: routing frequently interrupts the slowest cultivation attempts. This leads to an effective $2.4\times$-$6.7\times$ reduction in average cultivation time. These advantages are most pronounced in highly parallel circuits with substantial demand for magic states. Reducing average magic-state generation time is particularly valuable because the overall speed of a fault-tolerant quantum computer is governed by how quickly magic states can be produced and consumed.
@@ II-A Surface Codes
We model surface-code logical qubits as patches arranged on a two-dimensional grid [12]. This abstraction ensures that all required physical operations map cleanly onto a planar architecture. The edges of each patch play a central role: merging patches along specific edges performs joint parity measurements of the corresponding logical operators [12, 7]. Figure 2(b) shows a single surface-code patch with its $X$- and $Z$-type edges labeled. In our architecture, we employ double-qubit patches, represented as $2\times 1$ rectangles. Each such patch encodes two logical qubits within a single surface-code region. This layout places the $X$ and $Z$ operators for each qubit along the same boundary of the rectangle, enabling more efficient multi-qubit interactions and simplifying lattice-surgery operations (Figure 2(c)).
@@ II-B Pauli Basis Computation via Lattice Surgery
Input circuits are assumed to use the standard Clifford+T gate set [8, 19, 22]. Our compiler transforms Clifford+T circuits into sequences of Pauli product rotations by rewriting gates into their Pauli rotation equivalents (e.g., $S=Z_{\pi/4}$, $T=Z_{\pi/8}$). All Clifford operations are commuted to the end of the circuit and absorbed into measurements using the Clifford tableau method [1, 25].
@@
Lattice surgery [12] provides an efficient method to execute Pauli products through merge and split operations, as shown in Figure 3. These form the native instruction set of lattice surgery architectures.
@@ II-C Magic State Cultivation
Universal quantum computation requires magic states, which cannot be initialized fault-tolerantly. Traditional approaches rely on distillation, a process that consumes thousands of noisy resource states to produce a single high-fidelity one, leading to extremely large overheads [3, 8].
Recent work by Gidney et al. [9] introduced cultivation, a far more efficient approach using the same physical resources as a single logical qubit. Cultivation proceeds in three stages. First, injection creates the initial encoded T state in a low-distance code. Second, cultivation gradually increases reliability using “check-grow-stabilize” cycles with post-selection (shots are discarded when any error is detected). Finally, the escape stage quickly boosts the code distance to prevent the highly reliable state from degrading. At this point the code is “grafted” into a matchable surface-code-like patch. Cultivation’s efficiency enables more practical algorithm implementations and compact architectures [10], but existing scheduling algorithms have not thoroughly studied its implications for resource management.
@@ II-D Other approaches to FTQC with Surface Codes
Prior work on lattice-surgery compilation spans several distinct approaches. LaSsynth [27] introduces a SAT encoding of lattice surgery that can produce depth-optimal implementations of small-scale primitives. Although optimal, the method does not scale to the compilation of full application-level circuits. Building on SAT-based formulations, DASCOT [20] presents an encoding of the Lattice Surgery Scheduling Problem together with heuristic approximations to optimal solutions; however, it assumes that magic states are always available when needed and operates directly on Clifford+T circuits rather than translating them into Pauli products. Similarly, liblsqecc [17] proposes a framework for directly executing Clifford+T gates but considers only earlier generations of magic-state distillation and does not incorporate recent cultivation techniques. An alternative approach uses factory-less distillation layouts with static scheduling to exploit locality [11].
Architectural work has also explored ways of mitigating resource bottlenecks. LSQCA [14] introduces a design in which surface-code patches expected to be used soon are kept in sparse compute regions, while others are stored in denser regions; the movement between these regions helps conceal the latency of magic-state generation.
@@ III Scheduling
The goal of the scheduler is to solve the Lattice Surgery Scheduling Problem (LSSP). Given a circuit comprised of a sequence of Pauli products with $L$ qubits, the scheduling algorithm must compute an ordering of the Pauli products that minimizes volume, thereby minimizing the probability of error. The scheduler must also determine the grid layout of the logical qubits, with $L$ data qubits and additional ancilla qubits for routing and magic injection, for a total of $N>L$ qubits.
@@
Here $\epsilon$ is the error rate per physical qubit per cycle, $\epsilon_{th}$ is the error threshold, and $d$ is the code distance. Equation 1 assumes that physical errors (e.g. bit-flips, phase flips) are independent and identically distributed, both in time and space. In practice the error rate can be much more complex with dependent effects, governed by the underlying hardware. We follow other analyses in this area [8, 23] and use the simple error model. Minimizing volume minimizes errors under this error model.
@@ III-B Magic Injection Using Cultivation
We conducted Markov Chain Monte Carlo (MCMC) simulations using the survivorship probabilities of each stage of magic state cultivation outlined in Figure 15 of Gidney et. al. [9]. We simulated the cultivation of fault-distance-3 magic states so as to attain an error probability of $p_{L}\approx 4\times 10^{-6}$. Fitting the simulation data to Equation 3 yields a value of $\lambda=0.00227$. The duration of cultivation in scheduling cycles can be computed by taking samples from this exponential distribution then dividing by the code distance. Figure 7 shows the cumulative distribution function of $10^{6}$ simulated runs of cultivation.
The relationship between noise model, intermediate fault-distance, and final code distance is highly complex. Furthermore, recent work has demonstrated how cultivation can be made more time efficient  [28, 24]. For this reason, we match the parameters presented in Figure 15a of Gidney at. al. [9], but leave $\lambda$ as a controllable hyperparameter of our simulation framework.
@@ III-C Packing Steiner Forests
The scheduling algorithm attempts to maximize the number of non-overlapping, independent Pauli products scheduled in each cycle. As described previously, each Pauli product is scheduled as a tree on the grid. Ideally this is a Steiner tree, which connects all the terminals (the set of leaves representing the Pauli product operators) via the shortest paths. However, finding a Steiner tree is in general an NP-hard problem [18] so we find approximate Steiner trees using a variant of the shortest path heuristic [15]. This yields a tree with weight (path length) within 22/S2-2/S of the Steiner tree, where $S$ is the number of terminals. Although there are other algorithms with improved lower bounds on the approximation, in practice the shortest path heuristic provides sufficient accuracy for the sizes of trees for typical circuits.
Maximizing the number of simultaneously scheduled Pauli products is akin to solving the Steiner forest packing problem [16], which is also NP-hard. We find an approximate solution with a greedy MINFIT algorithm, as shown in Algorithm 1. The algorithm begins with the set of available Pauli products (those that have no parents in the task graph) and all squares in the grid. In each iteration of the while loop (line 4), for every Pauli product  an approximate Steiner tree is calculated for the available patches (line 6), and the Pauli product with the shorter path (the minimum fit) is chosen as the next to schedule (lines 7 to 10). That Pauli product is added to the Pauli products to schedule next (line 12), and removed from the list of available Pauli products (line 13), while the patches used by the tree for the Pauli product are removed from the available patches (line 14). Once all available Pauli products have been attempted, the cycle is complete. The MINFIT algorithm is repeated every cycle until no more unscheduled Pauli products remain, resulting in a schedule that indicates which Pauli products are to be executed at each cycle.
@@ IV Pure Magic Scheduling
One of the most significant performance constraints in surface-code quantum computation is the production of magic states. Even with cultivation, T-state preparation remains slow. Using the parameters reported in Gidney et. al. [9], our simulations indicate an average cultivation time of roughly 26 logical cycles. The latency of cultivation remains a large source of run time overhead, which is detrimental to performance for near-term architectures.
@@ V-A Test Circuits
We used a set of 17 benchmark circuits drawn from the Benchpress  [^2] The ”Heisenberg” benchmarks are the ”Square Heisenberg” in the Benchpress repository. quantum circuit suite [21], as shown in Table I. These circuits span many domains of interest (Quantum Machine Learning, Ising Models, Heisenberg Models, Convex Optimization). The circuits were lowered from their initial gate set to the standard Clifford+T gateset using BQSKit [2]. Clifford+T circuits were then converted to a sequence of Pauli products using the algorithm described in Silva et. al. [25]. The layers are derived from the task graph of the resulting Pauli products.
@@ V-D Comparison with Prior Work
This work is closest in nature to that of Silva et. al. [25]. In their paper, they present results with a set of specially generated benchmark circuits. For our work, we chose instead to use well-known benchmarks from the Benchpress suite. However, we have run our schedulers on the Silva et. al. benchmarks for comparison, using the versions they label “transpiled”. We present our scheduler performance compared to the results given in $E(N)$ column in Table 3 from Silva et. al. Specifically, we use our bus routing scheduler only, to get as close a comparison as possible to their results.
@@ VI-B Compatibility with other codes and execution models
Although we present the Pure Magic architecture in the context of the surface code, the conceptual approach is not code specific. Any scheme in which magic states may be cultivated locally (e.g. with color codes and other codes [30, 26]) could adopt a similar strategy. The key requirement is that ancillary space dedicated to routing be reallocated towards cultivation.
@@ VI-C Comparing and benchmarking QECCs
In general, QECCs are compared based on their qubit overhead, error suppression capabilities, and suitability for different hardware architectures. For example, while surface codes are robust, they are often criticized for their high physical to logical qubit encoding overhead. Whenever a code is compared against surface codes, the comparison [30, 29] uses static logical qubit layouts with dedicated bus and magic generation regions. This architecture has a fixed data to ancilla qubit ratio, and codes that use fewer physical qubits to encode a logical qubit tend to win against surface codes.